Given that 𝑇 = 3√ 32𝑐4βˆ’π‘ž /2 . Show that 𝑐 = 1/2 4√2𝑇3 + π‘ž Β Β 


1
Expert's answer
2021-11-04T20:09:15-0400

"𝑇 = \frac{\sqrt[3]{32𝑐^4−𝑞 }}{2}"


"2×T=\sqrt[3]{32𝑐^4-q}"


"2^3\u00d7T^3={32\ud835\udc50^4-q}"


"8T^3={32\ud835\udc50^4-q}"

"8T^3+q={32\ud835\udc50^4}"

"\sqrt[4]\frac{8T^3+q}{32}={𝑐}"


"c=\frac{1}{2}\sqrt[4]{2T^3+q}"